A black body is kept at a temperature of 2500K. If volume of body is reduced reversibly and adiabatically to 1/100 of its initial volume what will be the final temperature?
1
Expert's answer
2019-11-12T17:23:50-0500

The energy density per unit volume can be calculated according to Wien's radiation law


"u=\frac{8\pi h\nu^3}{c^3}e^{-h\nu/kT},"


and since the volume reduced adiabatically to 1/100 of the initial volume, the energy density increases 100 times, and while other constants ans variables remain the same, the black body acquired the new temperature:


"u_1=100u=\frac{8\pi h\nu_1^3}{c^3}e^{-h\nu_1/kT_1},"

"\frac{u_1}{u}=100=\frac{\nu_1^3}{\nu^3}e^{h(\nu/kT-\nu_1/kT_1)},\\
\space\\
100\frac{\nu^3}{\nu_1^3}=e^{h(\nu/kT-\nu_1/kT_1)}."

How to find the frequency of the radiation "\nu"? Wien's displacement law is what will help us:


"\lambda=\frac{c}{\nu}=\frac{b}{T},\\
\space\\
\nu=cT/b,\\
\space\\
100\frac{T^3}{T_1^3}=e^{h(c/kb-c/kb)}=e^0=1,\\
\space\\
T_1=T\sqrt[3]{100}=2500\sqrt[3]{100}=11604\text{ K}."

So hot!